# Mpemba effect on non-equilibrium active Markov chains

### v1

@@ Abstract
The qualitative reasoning of the occurrence of the Mpemba effect in the energy landscape with multiple local minima is given in terms of the effective trapping of the initially prepared colder system in the metastable states leading to its slower relaxation. In the context of the current model, a passive random walker initiated at a hot temperature does not experience any trap and explores the energy landscape much faster than the initially prepared colder system. On the contrary, an active random walker, be it initiated at a hot or cold temperature or even for a significant barrier height of the landscape, may not experience any metastable state for intermediate values of its activity or persistence as it can now explore the phase space with additional frequency characterized by the active rates. Such a phenomena suppresses the Mpemba effect in the presence of activity. However, the activity can also act otherwise. For example, the active random walker can experience momentary trap in the energy landscape due its persistence in a particular state leading to an activity induced Mpemba relaxation. We note that a similar phenomena was also observed in active colloidal systems diffusing in a potential landscape [29]. Naturally, the activity induced or suppressed behavior is a complex convolution of the initial temperature of the particle, barrier heights of the underlying energy landscape and the active, diffusive rates.
@@
Figure 5: Oscillatory Mpemba effect characterised by the zero crossings in $\Delta L_{1}(t):=L^{hot}_{1}(t)-L^{cold}_{1}(t)$ for two initially prepared hot and cold systems as they relax to a common steady state. We measure $\Delta L_{1}(t)$ as a function of time for varying flipping rate $\gamma$ which shows a pronounced effect of the activity. The oscillations tend to go down with increased $\gamma$ as the active walker becomes a diffusive one in this limit [55]. Here, we have set $\epsilon=9.0$ throughout the analysis.
@@
Oscillatory Mpemba effect: As shown in Fig. (2), the eigenvalues of the transition matrix $\boldsymbol{W}(T_{b})$ can be complex. As a result, both the time evolution of the probability distribution and the distance measures will have oscillatory terms. Figure 4 illustrates the time evolution of two different distance measures, namely $L_{1}$ and KL, for two systems $A$ and $B$ that are initiated at two different temperatures $T_{h}$ and $T_{c}$ respectively with $T_{h}>T_{c}$. We show that unlike the previous results of the Mpemba effect where the distance measures decay monotonically with time, here the distance measures show non-monotonic oscillatory-like decay across all the measures. As a result, the trajectories of the initially hot and cold system cross multiple times. This also indicates that the so-called $|a_{2}(T,T_{b})|$-criterion which is used for monotonic relaxations can not be a rational measure for non-equilibrium systems with complex eigenvalues.
@@
To have a better representation for the oscillations, we plot the difference of the $L_{1}$-norm for the hot and the cold system namely $\Delta L_{1}(t):=L^{hot}_{1}(t)-L^{cold}_{1}(t)$ as a function of time $t$ for different activity rates $\gamma$ keeping $\epsilon$ fixed as shown in Fig. 5. The plots shows that $\Delta L_{1}$ crosses the origin multiple times before it asymptotically relaxes to zero. The number of oscillations decreases as function of $\gamma$ for a fixed $\epsilon$ and thus, for the higher values of the activity $\gamma$, we can observe a reentrant diffusive like behavior in the relaxation of the trajectories. This is not surprising as in the large flipping rate $\gamma$ limit, the active walker behaves like a passive one as can also be seen from the effective diffusion constant [55, 57].
Figure 6: Multiple crossings in the relaxation trajectories can emerge even with real eigenspectrum of the active Markov chain (with $\gamma=\epsilon=1$). The result is independent of distance measures used as shown here with: (a) $L_{1}$-norm, and (b) Kullback-Leibler (KL) divergence. The results shown are for model parameters: $E_{1}=0.0$, $E_{2}=0.1$, $E_{3}=0.3$, $B_{12}=B_{21}=0.6$, $B_{13}=B_{31}=0.1$, $B_{23}=B_{32}=0.6$, whereas the temperatures of the hot, cold and the bath temperatures are chosen to be: $T_{h}=10.0$, $T_{c}=1.0$ and $T_{b}=0.1$, respectively.
@@
On the contrary, with real eigenspectrum, is it always ensured to have a single crossing of trajectories? We consider one such case but find that the distance functions can cross twice before they asymptotically merge with each other as illustrated in Fig. 6. Since there are even crossings, the hotter system lags behind the initially colder system leading to no net Mpemba effect in the large time. The same conclusion is drawn from the measure $|a_{2}(T,T_{b})|$ as well [see Fig. S2 in the SM]. Note that such a scenario is also observed in the non-equilibrium model of driven granular gases [58, 49] (albeit not universally but depending on the choice of the distance measure). Generically, this behavior can be attributed to the preparation of the systems in non-equilibrium conditions as is the case for the active systems here.
@@
Conclusions: In this paper, we have investigated the role of broken detailed balance in the anomalous relaxation phenomena. We unveil that the broken detailed balance conferred by the activity in a Markov chain system can showcase a myriad of unique relaxation phenomena such as the activity -induced and -suppressed Mpemba effect, and oscillatory Mpemba effect where relaxation trajectories cross multiple times. The latter is owing to the emergence of the complex eigenspectrum of the transition matrix in the presence of activity.
@@
Qualitatively, some of these phenomena can be reasoned with a heuristic description of ‘activity induced exploration on the energy landscape’. For instance, a cold walker can overcome metastabilities, kinetic traps or bottlenecks with the aid of active rates suppressing the Mpemba effect. However, activity can also act in favour of the Mpemba effect by trapping the walker of the initially colder system in a particular state with strong persistence. This would lead to the activity induced Mpemba effect. An intermediate persistence renders the walker to switch its state intermittently and thus, one would observe a non-monotonic relaxation behavior. Further analysis can shed light on this physical picture.
@@
While it is known that the active fluctuations can lead to an extensive variety of chemical and biological functions in living systems, its direct role in the quenching & relaxation dynamics pertaining to the Mpemba effect is yet to be characterized. Drawing inspiration from this simple analysis of an active Markov chain, we posit that some of these ideas maybe relevant to coarse-grained living systems such as fluctuating enzymes [59], proteins, or molecular machines [60] which continuously dissipate energy with the surrounding or exchange energy with the other molecules [53].
@@
It would be interesting to see whether experimental setups can be designed to capture such Mpemba like non-equilibrium phenomena from the transport & relaxation of bio-molecules through the engineered structures and cell biological systems [61] or even in simple active colloidal systems [62]. We believe that a synergy between the theory and experiment can provide a deeper conceptual understanding of this non-equilibrium phenomena.

### v2

@@ IV Activity harnessed Mpemba effect
The qualitative reasoning of the occurrence of the Mpemba effect in the energy landscape with multiple local minima is given in terms of the effective trapping of the initially prepared colder system in the metastable states leading to its slower relaxation. In the context of the current model, a passive random walker initiated at a hot temperature does not experience any trap and explores the energy landscape much faster than the initially prepared colder system. On the contrary, an active random walker, be it initiated at a hot or cold temperature or even for a significant barrier height of the landscape, may not experience any metastable state for intermediate values of its activity or persistence as it can now explore the phase space with additional frequency characterized by the active rates. Such realizations impede the Mpemba effect in the presence of activity. However, the activity can also act otherwise. For example, the active random walker can experience momentary trap in the energy landscape due its persistence in a particular state leading to an activity induced Mpemba relaxation. Summarizing, the activity induced behavior is a complex convolution of the initial temperatures; the active and diffusive rates of the particle; barrier heights of the underlying energy landscape and the activated escape rate from the metastable states [42, 43].
Figure 4: Inference of the Mpemba effect is independent of the choice of the distance measure. Illustration of the Mpemba effect using: (a) $L_{1}$-norm, and (b) Kullback-Leibler (KL) divergence. Here, the model parameters: $E_{1}=0$, $E_{2}=T_{b}$, $E_{3}=5T_{b}$, $B_{12}=B_{21}=6T_{b}$, $B_{13}=B_{31}=5T_{b}$, $B_{23}=B_{32}=4T_{b}$, $\gamma/R_{12}(T_{b})=593.65$, $\epsilon/R_{12}(T_{b})=148.41$, $T_{b}=0.1$ whereas the initially $hot$ and $cold$ temperatures are chosen to be: $T_{h}/T_{b}=30$ and $T_{c}/T_{b}=8$ respectively.
Figure 5: Illustration of the oscillatory Mpemba effect across different distance measures: (a) the $L_{1}$-norm, and (b) the Kullback-Leibler (KL) divergence for a system with broken detailed balance. The complex eigenspectrum can lead to oscillatory relaxations and distance measures. Here, the model parameters are: $E_{1}=0$, $E_{2}=T_{b}$, $E_{3}=3T_{b}$, $B_{12}=B_{21}=4T_{b}$, $B_{13}=B_{31}=5T_{b}$, $B_{23}=B_{32}=4T_{b}$, $\gamma/R_{12}(T_{b})=160.68$, $\epsilon/R_{12}(T_{b})=180.77$, $T_{b}=0.1$ whereas the temperatures of the hot and cold systems are chosen to be: $T_{h}/T_{b}=50$ and $T_{c}/T_{b}=5$ respectively.
@@
Figure 6: Oscillatory Mpemba effect characterized by the zero crossings in $\Delta L_{1}(t):=L^{hot}_{1}(t)-L^{cold}_{1}(t)$ for two initially prepared hot and cold systems as they relax to a common steady state. We measure $\Delta L_{1}(t)$ as a function of time for varying flipping rate $\gamma/R_{12}(T_{b})$ which shows a pronounced effect of the activity. The oscillations tend to go down with increased $\gamma/R_{12}(T_{b})$ as the active walker becomes a diffusive one in this limit [36]. Here, we have set $\epsilon/R_{12}(T_{b})=180.77$ throughout the analysis.
V Oscillatory Mpemba effect
@@
It is evident from Fig. 2 that the eigenvalues of the transition matrix $\boldsymbol{W}(T_{b})$ can be complex. As a result, both the time evolution of the probability distribution and the distance measures will have oscillatory terms. Figure 5 illustrates the time evolution of two different distance measures, namely $L_{1}$ and KL, for two systems $A$ and $B$ that are initiated at two different temperatures $T_{h}$ and $T_{c}$ respectively with $T_{h}>T_{c}$. We show that unlike the previous results of the Mpemba effect where the distance measures decay monotonically with time, here the distance measures show non-monotonic oscillatory-like decay across all the measures. As a result, the trajectories of the initially hot and cold system cross multiple times (hence referred as the Oscillatory Mpemba effect) and $|a_{2}(T,T_{b})|$-criterion is not suitable to depict such situation. Notably, the distance measures such as $L_{1}(t)$ and $KL(t)$ can contain superposition of multiple frequencies and not a unique one, which we discuss in detail in Sec. VI.
@@
To have a better representation for the oscillations, we plot the difference of the $L_{1}$-norm for the hot and the cold system namely $\Delta L_{1}(t):=L^{hot}_{1}(t)-L^{cold}_{1}(t)$ as a function of time $t$ for different activity rates $\gamma$ keeping $\epsilon$ fixed as shown in Fig. 6. The plots show that $\Delta L_{1}$ crosses the origin multiple times before it asymptotically relaxes to zero. The number of oscillations decreases as function of $\gamma$ for a fixed $\epsilon$ and thus, for the higher values of $\gamma$, we can observe a re-entrant diffusive like behavior in the relaxation of the trajectories. This is not surprising as in the large flipping rate $\gamma$ limit, the active walker behaves like a passive one as can also be seen from the effective diffusion constant [36, 44].
We delve deeper by analyzing the variation in active parameter space and the consequence to the oscillatory Mpemba effect. To this end, first we focus on two phase diagrams spanned by – (a) ratio of the active rates $\gamma/\epsilon$ and the temperature of the hot reservoir $T_{h}$, and (b) active rates $\epsilon$ and $\gamma$, keeping the other parameters fixed respectively (see Fig. 7). In the former case, the phase boundary remains insensitive to $T_{h}$ following an initial slow variation thus maintaining two distinct regions. In the latter case, the phase boundary changes if we vary $\epsilon$ and $\gamma$. It becomes evident that as $\gamma$ is increased, the domain of oscillations will gradually shrink indicating a vanishing Mpemba effect, as was indicated in Fig. 6. In addition, now we discuss few additional plots [see Fig. 8] that describes the behavior of the activity on the anomalous cooling but in the following planes of phase space: (a) variation of $\gamma$ with $T_{h}$, (b) variation of $\epsilon$ with $T_{h}$, and (c) simultaneous variation of $\gamma$, $\epsilon$ and $T_{h}$, while keeping every other parameter of the model that is not taken as a variation in the plane of the diagram is kept fixed. Figure 8(a) shows that for a fixed value of the model parameters, there exists a critical $\gamma$ that separates the region of oscillatory Mpemba with no Mpemba region, and with the increase in $\gamma$, the area of the oscillatory Mpemba region decreases as discussed earlier in Fig. 7. However, the behavior of the phase space variation is different in $\epsilon-T_{h}$ plane as shown in Fig. 8(b). Here, the critical value of $\epsilon$ that separates the region of oscillatory Mpemba with no Mpemba, remains the same for the chosen values of other parameters of the model. The combined analysis with the simultaneous variation of $\gamma$ and $\epsilon$ with $T_{h}$ is described in Fig. 8(c). However, the phase space now depicts the relevant timescales of the activity, i.e., $1/\gamma$ and $1/\epsilon$, that results in the oscillatory Mpemba effect as a function of the variation in $T_{h}$. These comprehensive phase diagrams provide a cue to asses the relevant active timescales in probing the oscillatory effect.
Figure 7: Phase diagram for the oscillatory Mpemba effect. Panel (a) illustrates the phase space region that shows the presence or absence of the Mpemba effect as the ratio of the activity $\gamma/\epsilon$ is varied along with $T_{h}$ keeping the other parameters fixed: $E_{1}=0$, $E_{2}=T_{b}$, $E_{3}=3T_{b}$, $B_{12}=B_{21}=4T_{b}$, $B_{13}=B_{31}=5T_{b}$, $B_{23}=B_{32}=4T_{b}$ and $T_{b}=0.1$. Panel (b) illustrates the variation in the phase space in the $\epsilon-\gamma$ plane setting $T_{h}/T_{b}=40$ and other parameters as in panel (a). The region for non-oscillatory Mpemba effect is seen to expand with large $\gamma$. Here, we have set $T_{c}/T_{b}=10$.
Figure 8: Phase diagram of the active Mpemba effect. Panel (a) illustrates the phase space region that shows the presence or absence of the Mpemba effect as the flipping rate $\gamma$ is varied along with $T_{h}$, the temperature of the hot system, whereas the other parameters are kept fixed at $E_{1}=0$, $E_{2}=T_{b}$, $E_{3}=3T_{b}$, $B_{12}=B_{21}=4T_{b}$, $B_{13}=B_{31}=5T_{b}$, $B_{23}=B_{32}=4T_{b}$, $\epsilon/R_{12}(T_{b})=40.17$ and $T_{b}=0.1$. Panel (b) illustrates the variation in the phase space in the $\epsilon$ versus $T_{h}$ plane but now keeping $\gamma/R_{12}(T_{b})=40.17$ fixed and the other parameters are same as in panel (a). Panel (c) illustrates the phase space of the oscillatory Mpemba effect in terms of relevant timescales of activity parameters: $1/\gamma$ and $1/\epsilon$ as a function of the variation in $T_{h}$. The other model parameters are kept fixed as per (a) and (b), and the temperature of the initially cold system in all the cases, are chosen to be $T_{c}/T_{b}=10$.
VI Presence of multiple frequency of oscillations in the oscillatory Mpemba effect
The oscillatory Mpemba effect or equivalently the oscillatory behavior in $L_{1}(t)$ (or $\Delta L(t)$) [see Fig. 5 and Fig. 6] can be presumed to have certain resemblance with that of a simple damped oscillator. In this section, we make an attempt to understand whether there is a similarity in the oscillatory behavior in $L_{1}(t)$ with that of a simple damped oscillator. For that let us expand the expression for the measure $L_{1}(t)$ in terms of eigenmodes: $L_{1}(t)=\sum_{i=1}^{3}|P_{i}(t)-\pi_{i}(T_{b})|$, where $P_{i}(t)$ denotes the occupation probability of site $i$ at time $t$ and $\pi_{i}(T_{b})$ denotes its final steady state probability at temperature $T_{b}$. The probabilities $P_{i}(t)$ for sites $i=1,2,3$ are obtained from the total probability vector $\boldsymbol{P}=(P_{1},P_{2},P_{3},P^{*}_{1},P^{*}_{2},P^{*}_{3})^{\mathbb{T}}$ [see Eq. (5)] by taking the first three components and their expansions in terms of eigenmodes can be written following the expansion of $\boldsymbol{P}$ [see Eq. (5)] as follows:
$P_{j}(t)=\pi_{j}(T_{b})+\sum^{6}_{i=2}a_{i}(T,T_{b})v_{ij}e^{\lambda_{i}t},\quad j=1,2,3,$ (7)
where $v_{ij}$ denotes the $j^{th}$ component of $i^{th}$ eigenvector $\boldsymbol{v}_{i}$ of the transition matrix $\boldsymbol{W}(T_{b})$ and $\lambda_{i}$ are its eigenvalues. For simplicity, while writing $P(t)$ from $\boldsymbol{P}(T,t)$ using Eq. (5), we have omitted $T$ from its parenthesis. Putting the expansion of Eq. (7) in the expression for $L_{1}(t)$, we obtain
$\begin{split}L_{1}(t)&=|\sum^{6}_{i=2}a_{i}(T,T_{b})v_{i1}e^{\lambda_{i}t}+\sum^{6}_{i=2}a_{i}(T,T_{b})v_{i2}e^{\lambda_{i}t}\\ &\ +\sum^{6}_{i=2}a_{i}(T,T_{b})v_{i3}e^{\lambda_{i}t}|=|\sum^{6}_{i=2}a_{i}(T,T_{b})e^{\lambda_{i}t}\sum^{3}_{j=1}v_{ij}|.\end{split}$ (8)
We recall that in the eigenspectrum of the transition matrix $\boldsymbol{W}(T_{b})$, the eigenvalue $\lambda_{1}=0$ corresponds to the steady state $\boldsymbol{\pi}(T_{b})$ and the remaining eigenvalues $\lambda_{i}$, $i=2,\ldots 6$ are in general complex with negative real parts. Since complex eigenvalues occur in pair, of the remaining eigenvalues there can be at most 4 complex eigenvalues, i.e., two pairs of complex numbers and in that case, the remaining one eigenvalue has to be real.
Consider this generic case where the eigenvalue $\lambda_{2}$ is real while $\lambda_{3,4}=\beta_{r}\pm\mathrm{i}\beta_{c}$ and $\lambda_{5,6}=\delta_{r}\pm\mathrm{i}\delta_{c}$ are complex eigenvalues. With this consideration and dropping the arguments of coefficients $a_{i}$ for simplicity, we can write $L_{1}(t)$ as
$\begin{split}L_{1}(t)&=|a_{2}e^{\lambda_{2}t}\sum^{3}_{j=1}v_{2j}+a_{3}e^{(\beta_{r}+\mathrm{i}\beta_{c})t}\sum^{3}_{j=1}v_{3j}+a_{4}e^{(\beta_{r}-\mathrm{i}\beta_{c})t}\sum^{3}_{j=1}v_{4j}+a_{5}e^{(\delta_{r}+\mathrm{i}\delta_{c})t}\sum^{3}_{j=1}v_{5j}+a_{6}e^{(\delta_{r}-\mathrm{i}\delta_{c})t}\sum^{3}_{j=1}v_{6j}|\\ &=|a_{2}\sum^{3}_{j=1}v_{2j}e^{\lambda_{2}t}+e^{\beta_{r}t}\big{[}(a_{3}\sum^{3}_{j=1}v_{3j}+a_{4}\sum^{3}_{j=1}v_{4j})\text{cos}(\beta_{c}t)+\mathrm{i}(a_{3}\sum^{3}_{j=1}v_{2j}-a_{4}\sum^{3}_{j=1}v_{2j})\text{sin}(\beta_{c}t)\big{]}\\ &\quad+e^{\delta_{r}t}[(a_{5}\sum^{3}_{j=1}v_{5j}+a_{6}\sum^{3}_{j=1}v_{6j})\text{cos}(\delta_{c}t)+\mathrm{i}(a_{5}\sum^{3}_{j=1}v_{5j}-a_{6}\sum^{3}_{j=1}v_{6j})\text{sin}(\delta_{c}t)]|\\ &=|\big{(}a_{2}\sum^{3}_{j=1}v_{2j}e^{\lambda_{2}t}+A_{\beta}e^{\beta_{r}t}\text{cos}(\beta_{c}t+\phi_{\beta})+A_{\delta}e^{\delta_{r}t}\text{cos}(\delta_{c}t+\phi_{\delta})\big{)}|,\end{split}$ (9)
where $A_{\beta}$, $A_{\delta}$ and $\phi_{\beta}$, $\phi_{\delta}$ denote the respective effective amplitudes and phases for the two modes of damped oscillations. Thus, in this case the oscillation in the quantity $L_{1}(t)$ is in effect a superposition of “two frequencies” corresponding to the imaginary part of the two pairs of complex eigenvalues.
On the other hand, a simple damped oscillator has a “single frequency” (say, $\nu_{0}$). Thus, in general there exists no one to one mapping between the oscillations in the quantity $L_{1}(t)$ [likewise the quantity $\Delta L(t)=L^{hot}_{1}(t)-L^{cold}_{1}(t)$] and a simple damped oscillator. Evidently, in certain cases, when there exists only one pair of complex eigenvalues, such a restrictive correspondence can be made. As an illustration, we recall the case considered in Fig. 5 – it should be noted that the eigenspectrum in this case has two pairs of complex eigenvalues and the same can be confirmed also from eigenvalue phase diagram Fig. 2(b). Evidently, there is no single frequency in this case as the preceding discussion suggests.
However, a particular case can be considered where $\beta_{r}$ happens to be the largest eigenvalue, in terms of its magnitude of its real part of the eigenvalue. In the large time limit, only the non-zero largest eigenvalue shall dominate and we can map the oscillatory decaying coefficient of the measure $\Delta L(t)$ [or $L_{1}(t)$] with the $a_{2}$ coefficient [see Eq. (30)] (which is also related with the largest non-zero eigenvalue) as
$a_{2}(T,T_{b})=\frac{A_{\beta}\text{cos}(\beta_{c}t+\phi_{\beta})}{\sum^{3}_{j=1}v_{2j}}.$ (10)
VII Multiple Mpemba effect even with real eigenspectrum
@@
Is it always ensured to have a single crossing of trajectories with real eigenspectrum? We consider one such case but find that the distance functions can cross twice before they asymptotically merge with each other as illustrated in Fig. 9. Since there are even crossings, the hotter system lags behind the initially colder system leading to no net Mpemba effect in the large time. The same conclusion is drawn from the measure $|a_{2}(T,T_{b})|$ as well [see Fig. 11]. Note that such a scenario is also observed in the non-equilibrium model of driven granular gases [41, 30] and molecular gases in the presence of non-linear drag [45] (albeit not universally but depending on the choice of the distance measure). Generically, this behavior can be attributed to the preparation of the systems in non-equilibrium conditions as is the case for the active systems here.
Figure 9: Multiple crossings in the relaxation trajectories can emerge even with real eigenspectrum of the active Markov chain (with $\gamma/R_{12}(T_{b})=\epsilon/R_{12}(T_{b})=148.41$). The result is independent of distance measures used as shown here with: (a) $L_{1}$-norm, and (b) Kullback-Leibler (KL) divergence. The results shown are for model parameters: $E_{1}=0$, $E_{2}=T_{b}$, $E_{3}=3T_{b}$, $B_{12}=B_{21}=6T_{b}$, $B_{13}=B_{31}=T_{b}$, $B_{23}=B_{32}=6T_{b}$, $T_{b}=0.1$, whereas the temperatures of the hot and cold systems are chosen to be: $T_{h}/T_{b}=100$ and $T_{c}/T_{b}=10$ respectively.
VIII Conclusion
@@
In this paper, we have investigated the role of broken detailed balance in the anomalous relaxation phenomena. We unveil that the broken detailed balance conferred by the activity in a Markov chain system can showcase a myriad of unique relaxation phenomena such as the activity -induced and -suppressed Mpemba effect, and oscillatory Mpemba effect where relaxation trajectories cross multiple times. The latter is owing to the emergence of the complex eigenspectrum of the transition matrix due to the breakdown of detailed balance in the presence of activity.
@@
Qualitatively, some of these phenomena can be reasoned with a heuristic description of ‘activity induced exploration on the energy landscape’. For instance, a cold walker can overcome metastabilities, kinetic traps or bottlenecks with the aid of active rates suppressing the Mpemba effect. However, activity can also act in favour of the Mpemba effect by trapping the walker of the initially colder system in a particular state with strong persistence. This would lead to the activity induced Mpemba effect. An intermediate persistence renders the walker to switch its state intermittently and thus, one would observe a non-monotonic relaxation behavior.
@@
While it is known that the active fluctuations can lead to an extensive variety of chemical and biological functions in living systems, its direct role in the quenching and relaxation dynamics pertaining to the Mpemba effect is not well understood. Drawing inspiration from this rigorous analysis of an active Markov chain, we posit that some of these ideas will be important to explore in the coarse-grained living systems such as fluctuating enzymes [39, 46], proteins, or molecular machines [47] that continuously dissipate energy with the surrounding or exchange energy with the other molecules [34].
@@
It would be interesting to see whether experimental setups can be designed to capture such Mpemba like non-equilibrium phenomena from the transport and relaxation of bio-molecules through the engineered structures and cell biological systems [48] or even in simple active colloidal systems [49]. We believe that a synergy between the theory and experiment can provide a deeper conceptual understanding of this non-equilibrium phenomena.
IX Acknowledgement